Which equation has a slope of -3 ?
A. 2y+6x=1
B. 2y-6x=1
C. 2x+6y=-1
D. 6x-2y=1

Answer by stanbon(57203)

Which equation has a slope of -3 ?
A. 2y+6x=1
slope = -3
-----
B. 2y-6x=1
slope = 3
-----
C. 2x+6y=-1
slope = -1/3
-----
D. 6x-2y=1
slope = 3
===============
Cheers,
stan H.